If a natural number n is of the form 3p+2, then n3 is also a number of same type varify.
Answer by venugopalramana(3286)

N=(3P+2)
N^3=(3P+2)^3=(3P+2)(3P+2)^2....SINCE N IS A NATURAL NUMBER P IS ALSO AN INTEGER.SO (3P+2)IS AN INTEGER.SO(3P+2)^2 IS AN INTEGER.SO
N^3=INTEGER*(3P+2)..SO IT IS OF THE FORM 3P+2